ailsa:

What is the percent by mass of potassium in K3Fe(CN)6?

Nnesha:

\[\large\rm Percent~by~Mass=\frac{Mass ~~of~ ~Solute}{Mass ~~of~~ Solution} \times 100\]

Nnesha:

\[\huge\rm \color{orange}{K}_\color{red}{3}Fe(CN)_\color{blue}{6}\] \[\rm Percent ~Mass ~of~K= \frac{\color{red}{3} \cdot 39.0983 }{Molar~Mass~of~the ~compount} \times 100\]

Nnesha:

\[\large\rm K= 3 \times 39.0983= ?\]\[\large\rm Fe= 1 \times 55.85= ?\]\[\large\rm C=6 \times 12.01=??\]\[\large\rm N= 6 \times 14.01=??\] add Em up. That will be the denominator MM

Nnesha:

I found this awesome pic PercentMassFromMM.png

sillybilly123:

fookin' awesome graphics!!

EclipsedStar:

Moved to the Chemistry section from Mathematics.

Vocaloid:

Closed due to answered.